You have $292,491 to invest in a stock portfolio (this amount is your original wealth). Your choices are Stock H, with an expected return of 14.42 percent, and Stock L, with an expected return of 10.48 percent. Legal constraints require you to invest at least $37,755 in stock L. If your goal is to create a portfolio with an expected return of 18.76 percent on your original wealth, what is the minimum amount you must borrow (and subsequently repay) at the risk free rate of 4.75 percent to achieve your goal? Answer in $ to two decimals. Thank you!!
Answer #1

Let the amount borrowed be x

Portfolio return is equal to weighted average return

37755*10.48% + (292491+x-37755)*14.42% - x*4.75% = 292491*18.76%

3956.724 + 36,732.9312 + 0.1442x – 0.0475x = 54,871.3116

0.0967x = 14,181.6564

X = $146,656.22

Hence, amount required to be borrowed = $146,656.22